OpenStudy (anonymous):

The length of a chord of a circle is 24 and its distance from the center is 5. What is the length of the diameter of the circle

OpenStudy (anonymous):

 Use Pythagorean THeorem c^2 = a^2 + b^2 c^2 = 5^2 + 12^2 c^2 = 25 + 144 c^2 = 169 c = 13

OpenStudy (anonymous):

THat's the radius. Double it to get the diameter. 13*2 = 26 THe diameter is 26

OpenStudy (anonymous):

what is the number above 24

OpenStudy (anonymous):

5

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

You're welcome :)